OpenStudy (kikuo):

Can someone explain in physics terms how velocity and power are related to acceleration?

OpenStudy (allen27):

wheres my medal

OpenStudy (kikuo):

@allen27 That doesn't exactly explain this. Lets say the the mass is at rest on a frictionless platform. The resultant force acting on the mass is zero and the initial velocity, u, of the body is also zero. Now a constant force, F, is applied to the mass and its start to move with a velocity. After the mass has moved a distance, s, away from its initial position, its velocity will be v m/s. Use the equation of motion to calculate the acceleration of the mass. v^2 = u^2 + 2as a = (v^2 - u^2)/2s Then , power = m x a x v 2nd method) P = m x g x v------(1) Replace a = f/m in (1) P = Fv

OpenStudy (welshfella):

power = work / time work = force * distance so power = force * distance / time = force * velocity P = F.v

OpenStudy (agent0smith):

More power means more acceleration - since the same amount of work (change in KE) can be done in less time